### v1

@@ 5 Prompt Generation, Evaluation and Control
This survey examines how an architecture that is good at System 1 tasks can be prompted to solve System 2 tasks. The Chain-of-thought paper showed how a simple command could prompt an LLM to perform reasoning steps, yielding much better performance in math word problems. Since then much research has further explored this approach, trying to build the ultimate general problem solver for System 1 and System 2 problems.
@@
Following the pipeline of Section 2.3, the prompts must (1) generate the reasoning steps, (2) evaluate the answer to the steps, and (3) control the number of steps that are generated, the shape (or complexity) of the reasoning process must be controlled. We will now briefly discuss the three stages. Please refer to Figure 1 for a diagram of the different approaches for the generation, evaluation, and control of reasoning steps, and to Table 1.222We show the approaches in the Figure in their main category only. Some approaches show innovations in two categories, and are shown twice. (Since all approaches have a generation, an evaluation, and a control aspect, all could in principle occur three times—all three columns can be found in Table 1).
Prompt for Step Generation
The first order of business is to create a prompt that instructs the LLM to generate reasoning steps. The problem must be split into substeps. This can be achieved with a problem-specific prompt that contains elements of the problem, such as: “First calculate how many marbles Mary had originally, then how many her friend had, and finally how many they had together.”
@@
In general, it is possible to prompt an LLM to fill in the blanks in a step-by-step fashion. In the papers that we will discuss, there are three main approaches for generating the step-by-step prompt. The prompt may be (1) handcrafted for the problem by the researchers (hand-written prompt), or (2) the prompt or prompts may come from an source that is external to the model, such as another model or a dataset (prompt using external knowledge), or (3) the model itself can be prompted to generate a (series of) prompt(s) to analyze the problem (model-generated prompt). As we will see, all three approaches have their advantages and disadvantages.
@@
After the prompt has been generated and the model has answered it, the next step in the reasoning pipeline is to evaluate the answer. Again, we see three main approaches for substep evaluation. First, the steps may be evaluated by (1) the model itself (self-assessment). Second, (2) an external program can be used to evaluate the steps. For example, when the steps are expressed as computer code, an external interpreter or compiler can be used to check the validity and the outcome (tool-based evaluation). Finally, (3) an external model can be used, LLM or otherwise. For example, in robotics, an external physics model can determine if certain actions are physically possible (external model validation).
Perform Control of Reasoning Steps
@@
A reasoning process that consists of multiple steps is a sequential decision process (Littman, 1996). When a single chain of reasoning steps is generated, the control flow of the reasoning process is simple: greedily evaluate the first step and then the next one, if present. The control flow of the reasoning process may also be more intricate. Some reasoning problems can be divided into multiple subproblems. To execute, evaluate and combine the results of all substeps, a separate controller may be needed. This controller can be a prompt or an external algorithm.
@@
Again, we distinguish three approaches. Most papers use (1) a greedy selection approach: a single prompt with a single chain of steps is generated, and these steps are directly executed and followed. The second approach (2) is to generate an ensemble strategy of reasoning steps, evaluate them, combine the individual results, and present them as the result of the ensemble. Finally, (3) a full tree-search or a reinforcement learning (RL) algorithm can be used as scaffolding. In this case, when a step is followed and evaluated, the LLM can roll back and try a different reasoning step. This is a breadth-first search approach (Plaat, 2020). Going further, a full reinforcement learning approach can be used (Sutton and Barto, 2018; Plaat, 2022) to find an optimal policy for the sequential decision process. A full Markov Decision Process of state, action, transition, and reward function is specified, and step control can become a process where prompts are generated dynamically.
| Approach | Domain | Step generation | Step evaluation | Step control |
| Scratchpad (Nye et al., 2021) | math word | hand-wr/supervised | - | greedy/1prompt |
| Chain-of-thought (Wei et al., 2022b) | math word | hand-written | - | greedy/1prompt |
| ZS-CoT (Kojima et al., 2022) | math word | hand-written | - | greedy/1prompt |
| Auto-CoT (Zhang et al., 2022) | math word | model-generated | - | clustering |
| Complexity (Fu et al., 2022) | math word | hand-written | self-consistency | greedy/1prompt |
| Self-ask (Press et al., 2022) | math word | external knowledge | LLM | multi-hop questions |
| Self-verification (Weng et al., 2022) | math word | hand-written | back-verify | ensemble |
| Self-consistency (Wang et al., 2022b) | math word | hand-written | majority | ensemble |
| Codex (Chen et al., 2021) | code | - | tool-based | - |
| Self-debugging (Chen et al., 2023) | code | hand-written | tool-based | greedy |
| Fun-search (Romera-Paredes et al., 2024) | code | hand-written | tool-based | evolutionary algorithm |
| LLaMEa (van Stein and Bäck, 2024) | code | hand-written | tool-based | evolutionary algorithm |
| MathPrompter (Imani et al., 2023) | math | hand-written | tool-based | ensemble |
| Program-of-thoughts (Chen et al., 2022) | math word | hand-written, Codex | Python+Consist. | decouple reason/compute |
| Program-aided-language (Gao et al., 2023) | math word | hand-written, Codex | NLP/Python | ensemble |
| Refiner (Paul et al., 2023) | math word | finetune | critic model | gen/crit feedback |
| Self-corrector (Welleck et al., 2022) | math word | finetune | corrector model | gen/corr feedback |
| Self-improvement (Huang et al., 2022a) | math word | finetune | self-assessment | CoT/consistency |
| Say-can (Ahn et al., 2022) | robot | model-generated | external model | greedy |
| Inner-monologue (Huang et al., 2022b) | robot | hand-written | various | greedy |
| Self-taught-reasoner (Zelikman et al., 2022) | math word | finetune | augmentation | greedy/feedback |
| Least-to-most (Zhou et al., 2022) | math word | hand-written | self-assessment | curriculum |
| Progressive-hint (Zheng et al., 2023) | math word | model-generated | self-assessment | stable prompt |
| Self-refine (Madaan et al., 2023) | math word | model-generated | self-assessment | greedy/feedback |
| Tree-of-thoughts (Yao et al., 2024) | puzzles | model-generated | self-assessment | BFS/DFS |
| Buffer-of-thoughts (Yang et al., 2024) | math word | thought template | self-assessment | buffer manager |
| Beam-search (Xie et al., 2024) | math word | model-generated | self-assessment | Beam Search |
| ReAct (Yao et al., 2022) | action | external knowledge | self-assessment | reinforcement learning |
| Reflexion (Shinn et al., 2024) | decision | model-generated | ext model | reinforcement learning |
| Voyager (Wang et al., 2023) | Minecraft | model-generated | Minecraft | reinforcement learning |
Domain
Many papers are applied to math word problems (natural language descriptions of math problems). Math problems were the original inspiration for the experiments with reasoning in LLMs. Other application domains include autonomous agents, robotic movement, generating computer programs, and playing computer games. We will discuss these in more detail with the individual approaches.
@@
Table 1 lists the papers of this survey. They are listed by the domain they work on, the type of prompt generation, the evaluation of the result, and the control method. The approaches in the table are grouped, divided by horizontal lines.
The first group, from Scratchpad to Self-ask, focuses on creating a prompt that generates the reasoning steps. The entries in the cells of this column are shown in bold, highlighting the focus of the approaches. The approaches in this group can be considered to be the start of the field of LLM-reasoning. The Chain-of-thought approach is especially an inspiration for many works. The prompts are often written “manually” by the researchers, the steps are encoded in one prompt, and step control is greedy. There is no specific evaluation of the steps, other than comparing results to the benchmark. The Scratchpad approach is special in that it uses supervised learning, not prompt-learning; the work showed that LLMs can be made to generate internal reasoning steps by supervised learning, paving the way for the later prompt-based papers.
The second group, from Self-verification to Self-taught-reasoner, focuses on evaluation of the reasoning steps in the prompt. This column is shown in bold in the table. The approaches in this group aim to improve the Chain-of-thought results by reducing the error accumulation that occurs when multiple steps are taken in a reasoning chain. A variety of step control methods is used by these approaches, which is discussed in more detail later. Note that not all approaches use natural language problems (often math word problems). For example, the subgroup of Codex to Program-aided-language focuses on formal languages. They generate code or math equations, typically in Python, to formalize the steps of the reasoning problem, or as result of the task. LLMs are quite good at code generation, and these approaches typically achieve good performance. The use of code also allows the approaches to call external programs such as interpreters and debuggers to evaluate the correctness of the reasoning steps that are generated.
There is also a special subgroup, Refiner to Self-improvement, that does not use prompt learning but finetuning. Here, new data is generated based on reasoning exemplars, which is then used to further train the model. The extra data is often generated as a separate dataset, sometimes called critic or corrector.
@@
The third group, Least-to-most to Voyager, addresses step control (approaches shown in bold in this column). Whereas in the previous approaches the reasoning steps are written in a single, static, prompt, these approaches generate the steps in multiple, dynamic, prompts. This allows control of the space of reasoning steps. Various search control approaches are used, all in the form of an external algorithm that performs calls to the LLM with different prompts. The control methods range from simple greedy and depth-first search to elaborate beam search and reinforcement learning schemes.
@@
In summary, we see a diverse array of methods that often achieve high performance in reasoning about their respective domains. To better understand the approaches, let us discuss the techniques in more detail, starting with the generation of steps.
5.1 Generation of Steps
@@
Originally, LLMs performed poorly on math word problems (GSM8K (Cobbe et al., 2021)). Some different approaches were tried, for example scaling up the size of the LLM (Rae et al., 2021). The LLM architecture, based on transformers, is designed to produce a single token. When we prompt such an architecture to produce an answer, it does so. What we should do is prompt it to follow intermediate steps, answer those, and thus work towards the final answer, just as a student is taught to break down a complex problem into smaller steps. We should take the model by its hand and teach it to write down the intermediate steps, and combine the intermediate results (Nye et al., 2021).
This idea was used by Nye et al. (2021) in Scratchpads, a transformer model that performs multi-step computations by asking it to emit intermediate computation steps into a scratchpad. They train the model by supervised learning (not prompt-based in-context learning). Figure 2 shows an example. On experiments with addition, polynomial evaluation, and Python code execution, versions that produced the intermediate steps on a scratchpad performed considerably better than versions that did not.
@@
If supervised learning can produce intermediate steps, would prompt-learning be able to do so too?
5.1.1 Hand-written Prompt
@@
This question was studied by Wei et al. (2022b), amongst others. A basic way to instruct an LLM to generate steps by prompt-learning is to manually write a prompt for the large language model to follow the reasoning steps. They showed in their Chain-of-thought paper that with such a prompt the LLM follows such intermediate steps. When the LLM is prompted to rephrase information from the question as intermediate reasoning steps in its answer, the LLM performed much better than when it was prompted to answer a math problem directly, without reproducing the information from the question in its answer. The example from the Chain-of-thought paper is shown in Figure 3 Wei et al. (2022b). Performance figures were given in Section 3 on benchmarks.
@@
The substantial performance improvement by Chain-of-thought has caused much excitement and has opened up further research on reasoning with LLMs. In the original Chain-of-thought paper the prompts were handwritten by the researchers for the individual types of problems, and evaluations are conducted with five different benchmarks (not by an LLM).333The Chain-of-thought idea is about prompt generation, not about the evaluation or the search control of the reasoning steps. Hence, in Table 1 Chain-of-thought is labeled as greedy without an evaluation. In a later work the prompts were generated automatically by the LLM (Zhang et al., 2022).
@@
Kojima et al. (2022) go a step further. They show that the simple addition of a single text to the prompt (Let’s think step by step) significantly improves performance. Since this text does not contain problem-related elements, this can be considered as a form of zero-shot learning. Figure 4 compares the approaches. Experiments further show that with this addition to the prompt, significant performance gains are achieved on a diverse set of reasoning benchmarks, including arithmetic, symbolic, and logical reasoning.
5.1.2 Prompt using External Knowledge
Chain-of-thought shows that an LLM gives better answers to complex problems when it is guided to take individual steps. Prompts are written manually, from scratch, by the researchers.

### v2

@@ 3. Step Generation, Evaluation and Control
This survey examines how LLMs based on the transformer architecture can be prompted to solve multi-step reasoning tasks. The Chain-of-thought paper shows how a simple command could prompt an LLM to perform reasoning steps, yielding much better performance in math word problems. Since then much research has further explored this approach, trying to build stronger general problem solvers for other types of reasoning problems.
A typical approach to solve a complex problem is to subdivide it into smaller steps and to solve those. This approach is related to classical divide and conquer (Bellman, 1966). It consists of three stages. New steps are (1) generated, (2) evaluated, and the search of the generated steps is (3) controlled in some way. The in-context reasoning approaches that we survey follow a general three-stage pipeline (Madaan et al., 2023):
Generate: prompt the model for the generation of steps,
Evaluate: prompt for evaluation of the generated steps,
Control: prompt for control of the number of steps that are generated and how deep ahead the reasoning process will look.
@@
This three-stage pipeline is the basis of our taxonomy. We will now discuss the three stages more deeply; for ease of reference they are numbered according to the Subsection in which they are described in more detail (3.1, 3.2, 3.3). Please also refer to Figure 1, or Table 2, for a diagram of the categories and subcategories of different approaches for the generation, evaluation, and control of reasoning steps.111We show the approaches in the Figure in their main category only. Some approaches show innovations in two categories, and are shown twice. (Since all approaches have a generation, an evaluation, and a control aspect, all could in principle occur three times, and all three columns can be found in Table 2).
(3.1) Generation
@@
The first stage is to create a prompt that instructs the LLM to generate reasoning steps. The problem must be split into substeps. This can be achieved with a problem-specific prompt that contains elements of the problem, such as: “First calculate how many marbles Mary had originally, then how many her friend had, and finally how many they had together.” In general, it is possible to prompt an LLM to fill in the blanks in a step-by-step fashion. In the papers that we discuss, there are three main approaches for generating the step-by-step prompt, numbered with the Subsection in which the approaches are described. First, the prompt may be handcrafted for the problem by the researchers: (3.1.1) hand-written prompt. Second, the prompt or prompts may come from a source that is external to the model, such as another model or dataset: (3.1.2) external knowledge-based prompt. Third, the model itself can be prompted to generate a (series of) prompt(s) to analyze the problem (3.1.3) model-generated prompt. As we will see, all three approaches have their advantages and disadvantages.
@@
After the prompt has been generated and the model has answered it, the next stage is to evaluate the quality of this answer. Such evaluation is often necessary to improve and perform well on the benchmark. Again, we see three main approaches for substep evaluation. First, the steps may be evaluated by the model itself: (3.2.1) self-assessment. Second, an external program can be used to evaluate the steps. For example, when the steps are expressed as computer code, an external interpreter or compiler can be used to check the validity and the outcome: (3.2.2) tool-based evaluation. Finally, an external model can be used, LLM or otherwise. For example, in robotics, an external physics model can determine if certain actions are physically possible: (3.2.3) external model validation.
(3.3) Control
@@
The third stage is control. A reasoning process that consists of multiple steps is a sequential decision process (Littman, 1996). When a single chain of reasoning steps is generated, the control flow of the reasoning process is simple: greedily evaluate the first step and then the next one, if present. The control flow of the reasoning process may also be more intricate. Some reasoning problems can be divided into multiple subproblems. To execute, evaluate and combine the results of all substeps, a separate controller may be needed. This controller can be a prompt or an external algorithm.
@@
Again, we distinguish three approaches. Most papers use a (3.3.1) greedy selection approach: a single prompt with a single chain of steps is generated, and these steps are directly executed and followed. The second approach is to generate an (3.3.2) ensemble strategy of reasoning steps, evaluate them, combine the individual results, and present them as the result of the ensemble. Finally, a full tree-search or a (3.3.3) reinforcement learning (RL) algorithm can be used as scaffolding. In this case, when a step is followed and evaluated, the LLM can roll back and try a different reasoning step (Plaat, 2020). Going further, a full reinforcement learning approach can be used (Sutton and Barto, 2018; Plaat, 2022) to find an optimal policy for the sequential decision process. In general a Markov Decision Process of state, action, transition, and reward function can be specified, and step control can become a process where prompts are generated dynamically under the control of an external RL algorithm, or as in-context reinforcement learning (ICRL) (Lee et al., 2023).
Approach Domain (3.1) Step generation (3.2) Step evaluation (3.3) Step control Result Scratchpad (Nye et al., 2022) math word hand-wr/superv - greedy/1prompt PolyEval +19%, Python +21% Chain-of-thought (Wei et al., 2022b) math word hand-written - greedy/1prompt GSM8K +39%, SVAMP +10%, ASDiv +2%, AQuA +11%, MAWPS +14%, CSQA +1.8%, StrategyQA +0.2% ZS-CoT (Kojima et al., 2022) math word hand-written - greedy/1prompt MultiArith =89%, GSM8K =70% Auto-CoT (Zhang et al., 2023) math word model-generated - clustering MultiArith +0.3%, GSM8K +1%, AddSub +3.5%, AQuA +0.7%, SingleEq +0.4%, SVAMP +0.6%, CSQA +1%, StrategyQA +0%, Letter +0.7%, Coin +2.7% Complexity (Fu et al., 2022) math word hand-written self-consistency greedy/1prompt GSM8K +7%, MultiArith +3%, Penguins +3% Self-ask (Press et al., 2023) math word ext knowledge LLM multi-hop q. Bamboogle =60%, 2Wiki =40%, Musique =15% Self-verification (Weng et al., 2023) math word hand-written back-verify ensemble GSM8K +4%, SingleEq +2%, AddSub +4%, MultiArith +3%, AQuA +3%, SVAMP +1% Self-consistency (Wang et al., 2023b) math word hand-written majority ensemble GSM8K +18%, SVAMP +11%, AQuA +12%, StrategyQA +6%, ARC-c +4% Codex (Chen et al., 2021) code - tool-based - HumanEval =70% Self-debugging (Chen et al., 2024) code hand-written tool-based greedy Spider +9%, MBPP +12%, TransCoder +12% Fun-search (Romera-Paredes et al., 2024) code hand-written tool-based evolutionary alg cap set 8 =512 LLaMEa (van Stein and Bäck, 2024) code hand-written tool-based evolutionary alg BBOB +11% MathPrompter (Imani et al., 2023) math hand-written tool-based ensemble MultiArith =92% Program-of-thoughts (Chen et al., 2023) math word hand-wr, Codex Python+Consist. split reason/cmput GSM8K =71%, SVAMP =85%, ASDIV =85%, AddSub =92%, MultiArith = 99% Program-aided-lang (Gao et al., 2023) math word hand-wr, Codex NLP/Python ensemble GSM8K =72%, SVAMP =79%, ASDIV =79%, SingleEQ =96%, SingleOP =94%, AddSub = 92%, MultiArith = 99%, Penguins = 93% Refiner (Paul et al., 2024a) math word finetune critic model gen/crit feedback SVAMP =72%, GSM8K =78% Self-correction (Welleck et al., 2023) math word finetune corrector model gen/corr feedback MathProgSynth =24%, LexConstrGen =98%, ToxicityControl =0.0% Self-improvement (Huang et al., 2023) math word finetune self-assessment CoT/consistency GSM8K =82%, DROP =83%, ARC-c =90%, OpenBookQA =94%, ANLI-A3 =68% Say-can (Ahn et al., 2022) robot model-generated external model greedy Kitchen =31% Inner-monologue (Huang et al., 2022) robot hand-written various greedy TableTop =90%, Kitchen =60% Self-taught-reasoner (Zelikman et al., 2022) math word finetune augmentation greedy/feedback CommonsenseQA =72% Least-to-most (Zhou et al., 2023) math word hand-written self-assessment curriculum SCAN =99% Progressive-hint (Zheng et al., 2023) math word model-generated self-assessment stable prompt AddSub +2%, MultiArith +0%, SingleEQ +2%, SVAMP +3%, GSM8K +5%, AQuA +1% Self-refine (Madaan et al., 2023) math word model-generated self-assessment greedy/feedback Sentiment +32%, Dialog +49%, CodeOptim +8%, CodeRead +28%, MathReason +0%, AcronymGen +25%, ConstrainedGen +30% Tree-of-thoughts (Yao et al., 2024) puzzles model-generated self-assessment BFS/DFS Game24 =74%, CreativeWriting , Crossword Buffer-of-thoughts (Yang et al., 2024) math word thought template self-assessment buffer manager Game24 +11%, GeoShapes +20%, Checkmate +51% Algorithm-of-thoughts (Sel et al., 2023) puzzles model-generated self-assessment in-context RL GSM8K =89%, StrategyQA =84%, Crossword Beam-search (Xie et al., 2024) math word model-generated self-assessment Beam Search GSM8K +6%, AQuA +9%, StrategyQA +5% ReAct (Yao et al., 2023) action external knowledge self-assessment reinf learning ALFWorld =34%, WebShop =10% Reflexion (Shinn et al., 2024) decision model-generated ext model reinf learning HumanEval =91% Voyager (Wang et al., 2024a) Minecraft model-generated Minecraft reinf learning 15 x faster
@@
Table 2 lists the main papers of this survey. We show the domain they work on, the type of prompt generation, the evaluation of the result, and the control method. These three categories of approaches—indicated by their Sections (3.1) generation, (3.2) evaluation, (3.3) control—are shown in the table as groups divided by horizontal lines. The first group in the Table, from Scratchpad to Self-ask, focuses on creating a prompt that generates the reasoning steps. The entries in the cells of this column are shown in bold, highlighting the focus of the approaches. The approaches in this group are the start of the field of LLM-reasoning. The Chain-of-thought approach is especially an inspiration for many works. The prompts are often written manually by the researchers for each problem; the steps are encoded in one prompt, and step control is greedy. There is no specific evaluation of the steps, other than comparing results to the benchmark. The Scratchpad approach is special in that it uses supervised learning, not prompt-learning; the work showed that LLMs can generate internal reasoning steps by supervised learning, paving the way for in-context works.
The second group, from Self-verification to Self-taught-reasoner, focuses on evaluation of the reasoning steps in the prompt. This column is shown in bold in the table. The approaches in this group aim to improve the Chain-of-thought results by reducing error accumulation that occurs when multiple steps are taken in a reasoning chain. A variety of step control methods is used by these approaches, which is discussed in more detail later. Note that not all approaches use natural language problems. For example, the subgroup of Codex to Program-aided-language focuses on formal languages. They generate code or math equations, typically in Python, to formalize the steps of the reasoning problem, or as result of the task. LLMs are quite good at code generation (Chen et al., 2021), and these approaches typically achieve good performance. The use of code also allows the approaches to call external programs such as interpreters and debuggers to evaluate the correctness of the reasoning steps that are generated.
There is also a special subgroup, Refiner to Self-improvement, that uses finetuning in addition to prompt learning. Here, new data is generated based on reasoning exemplars, which is then used to further train the model. The extra data is often generated as a separate dataset, sometimes called critic or corrector.
@@
The third group, Least-to-most to Voyager, addresses step control (shown in bold in this column). Whereas in the previous approaches the reasoning steps are written in a single, static, prompt, these approaches generate the steps in multiple, dynamic, prompts. This allows control of the space of reasoning steps. Various search control approaches are used, all in the form of an external algorithm that performs calls to the LLM with different prompts. The control methods range from simple greedy and depth-first search to elaborate beam search and reinforcement learning schemes.
The last column of the Table summarizes reported benchmark results. The ’=’ symbol indicates absolute scores on the benchmarks, while ’+’ indicates relative improvement in percentage points over a baseline LLM, typically GPT-3.5. Results vary strongly, both between approaches and within a single approach between benchmarks. Also, different LLMs were used, from early stage to more mature models, open and commercial, and the baselines differ. For some benchmarks, such as the Creative Writing benchmark in Tree-of-thoughts, results are best reported qualitatively. The source papers provide more measurement details, Section 4.1 discusses when to which approach to use for different applications.
@@
In conclusion, we see a diverse array of methods that often achieve high performance in reasoning on their respective domains. To better understand the approaches, we discuss them in more detail, starting with the generation of steps.
3.1. Generation of Steps
@@
Originally, LLMs performed poorly on math word problems such as GSM8K (Cobbe et al., 2021). Different approaches were tried unsuccessfully, for example scaling up the size of the LLM (Rae et al., 2021). The LLM architecture, based on transformers, is designed to produce a single token. When we prompt such an architecture to produce an answer, it does so. What we should do instead, is to prompt it to follow intermediate steps, answer those, and thus work towards the final answer, just as a student is taught to break down a complex problem into smaller steps. We should guide the model to explicitly produce intermediate steps, and combine the intermediate results. This idea was used by Nye et al. (2022) in Scratchpads, a transformer model that performs multi-step computations by asking it to emit intermediate computation steps into a scratchpad. They train the model by supervised learning (not prompt-based in-context learning). Figure 2 shows an example. On experiments with addition, polynomial evaluation, and Python code execution, versions that produced the intermediate steps on a scratchpad performed considerably better than versions that did not, going from 35% to 95%, from 32% to 51%, and from 30% to 42% accuracy, respectively.
@@
If supervised learning can produce intermediate steps, would prompt learning be able to do so too?
3.1.1. Hand-written Prompt
@@
This question was studied by Wei et al. (2022b), amongst others. A basic way to instruct an LLM to generate steps by prompt-learning is to manually write a prompt for the large language model to follow the reasoning steps. When the LLM is prompted to rephrase information from the question as intermediate reasoning steps in its answer, the LLM performed much better than when it was prompted to answer a math problem directly, without reproducing information from the question in its answer in multiple steps. The example from the Chain-of-thought paper is shown in Figure 3. Table 1 shows that the largest accuracy increase is on GSM8K, from 16% to 47%.
@@
The idea that an LLM can be made to follow step-by-step instructions, and the performance improvement by Chain-of-thought have caused much excitement and have opened up further research on reasoning with LLMs. In the original paper the prompts were handwritten by the researchers for the individual types of problems, and evaluations are conducted with benchmarks (not by an LLM).222The Chain-of-thought idea is about prompt generation, not about the evaluation or the search control of the reasoning steps. Hence, in Table 2 Chain-of-thought is labeled as greedy without an evaluation. In a later work the prompts were generated automatically by the LLM (Zhang et al., 2023), and evaluated.
@@
Kojima et al. (2022) go a step further. They show that the addition of a single text to the prompt (Let’s think step by step) significantly improves performance. Since this text does not contain problem-related elements, it is as a form of zero-shot learning. Figure 3 (third column) compares the approaches. Experiments further show that with this addition to the prompt significant performance gains are also achieved on other reasoning benchmarks, including arithmetic, symbolic, and logical reasoning (achieving 70% accuracy on GSM8K/PaLM when Self-consistency is also included).
3.1.2. Prompt using External Knowledge